# Zero-Shot Clinical Trial Patient Matching with LLMs

### v1

@@ 6 Discussion
In this work, we developed a zero-shot LLM-based pipeline to match patients to clinical trials. We achieved state-of-the-art performance on the 2018 n2c2 cohort selection challenge, achieving gains of +6 Macro-F1 and +2 Micro-F1 points over the prior best model without any feature engineering. Unlike prior methods, our system also provides interpretable rationale for its predictions, thereby allowing for human oversight and collaboration. Clinician evaluators disagreed with the LLM-generated rationales only 3% of the time when the LLM made a correct eligibility decision, and 25% of the time when it made an incorrect decision. Finally, we show that a preliminary retrieval step can preserve high performance while reducing token usage, and that thoughtful prompting strategies can lead to large cost savings.
@@
In contrast to prior work, we do no fine-tuning or updates to the underlying LLM’s weights. Nor do we provide in-context few shot examples. This illustrates the unique scalability advantages that LLMs such as GPT-4 offer, as they can be seamlessly repurposed for arbitrary trials so long as definitions for the criteria are known. In contrast to prior rule-based systems, the LLMs we test demonstrate robust clinical language understanding and general reasoning capabilities.
@@
If our results generalize, then LLMs represent a viable solution for clinical trial patient identification given their accuracy, efficiency, and cost-effectiveness. Our GPT-4-based system is accurate, as shown in Table 1, and is able to output coherent explanations for the majority of its decisions, as depicted in Figure 3. It is an order of magnitude faster than today’s clinical trials operate, as it was able to evaluate all 86 patients in roughly two hours versus the one patient per hour expected for Phase III trials (Penberthy et al., 2012). Our pipeline is also more cost-effective, as Table 3 shows that the cost for using GPT-4 with the ACIN strategy to screen a single patient is roughly \$1.55, whereas prior estimates for a Phase III trial peg the cost per patient at \$34.75111See Table 3, $38193.70/1099=\mathdollar 34.75$ cost per patient screened. (Penberthy et al., 2012), a gap that will likely widen as models become more efficient.
@@
There are major limitations of our work. First, the size of our dataset is dwarfed by the scale of most health systems’ EHRs, which underscores the need for using preliminary retrieval-based pipelines which can reduce the amount of tokens that an LLM must process to match each patient. Second, the benchmark we use – the 2018 n2c2 cohort challenge – is a significant simplification of actual trial matching. It only covers inclusion criteria, only includes English, and its formatting might not generalize to other EHRs. Third, although our LLM-based system achieved state-of-the-art results on this benchmark, further evaluations are required to asses the safety and failure modes of applying these models in production settings. Fourth, the use of proprietary models such as GPT-4 may not be feasible for health systems that do not have HIPAA-compliant Azure instances.
At a higher level, there are also open questions as to whether low enrollment in trials is a function of not being able to match patients (which this work addresses) or is instead primarily due to the stringency of the eligibility criteria themselves (Liu et al., 2021). While beyond the scope of this work, we believe that improving the design of eligibility criteria presents another interesting research direction which could benefit from the application of more sophisticated models such as LLMs (Liu et al., 2021).

### v2

@@ 6 Discussion
In this work, we developed a zero-shot LLM-based pipeline to match patients to clinical trials. We achieved state-of-the-art performance on the 2018 n2c2 cohort selection challenge, achieving gains of +6 Macro-F1 and +2 Micro-F1 points over the prior best model without any feature engineering. Unlike prior methods, our system also provides interpretable rationale for its predictions, thereby allowing for human oversight and collaboration. Clinician evaluators disagreed with the LLM-generated rationales only 3% of the time when the LLM made a correct eligibility decision, and 25% of the time when it made an incorrect decision. Finally, we show that a preliminary retrieval step can preserve high performance while reducing token usage, and that thoughtful prompting can lead to large cost savings.
@@
In contrast to prior work, we do no fine-tuning or updates to the underlying LLM’s weights. Nor do we provide in-context few-shot examples. This illustrates the unique scalability advantages that LLMs such as GPT-4 offer to health systems, as they can be seamlessly repurposed for arbitrary trials so long as definitions for the criteria are known. Stanford, for example, has over 250 active clinical trials in oncology alone. Thus, even if a few-shot or fine-tuned system only required 5-10 hand-labeled patients per trial, deploying such an approach would require 1,000’s of labels, plus constant updating with each new trial. This is avoided with a zero-shot approach.
@@
If our results generalize, then LLMs represent a viable solution for clinical trial patient identification given their accuracy, efficiency, and cost-effectiveness. Our GPT-4-based system is accurate, as shown in Table LABEL:tab:zero_shot, and is able to output coherent explanations for the majority of its decisions, as depicted in Appendix Figure LABEL:fig:clinican_rationale. It is an order of magnitude faster than today’s clinical trials operate, as it was able to evaluate all 86 patients in roughly two hours versus the one patient per hour expected for Phase III trials (Penberthy et al., 2012). Our pipeline is also more cost-effective, as Table LABEL:tab:prompt_strat shows that the cost for using GPT-4 with the ACIN strategy to screen a single patient is roughly \$1.55, whereas prior estimates for a Phase III trial peg the cost per patient at \$34.75111See Table 3, \$$38193.70/1099=\$34.75$ cost per patient screened. (Penberthy et al., 2012), a gap that will likely widen as models become more efficient.
Relying on LLMs poses risks as they have a propensity to ”hallucinate”, i.e. make up facts (Umapathi et al., 2023). Thus, we envision such an LLM-based trial matching system being deployed primarily as a ”pre-screener” that automatically flags all patients in an EHR across all possible trials, thereby allowing clinical research coordinators to focus only on patients with strong enrollment potential. This minimizes the harm of misclassifying a patient by providing a natural means for human-in-the-loop validation.
Limitations
@@
There are major limitations of our work. First, the size of our dataset is dwarfed by the scale of most health systems’ EHRs, which underscores the need for using preliminary retrieval-based pipelines which can reduce the amount of tokens that an LLM must process to match each patient. Second, while the benchmark we use – the 2018 n2c2 cohort challenge – is the most realistic and largest public clinical trial matching dataset currently available, it is still a significant simplification of actual trial matching. It only covers inclusion criteria, only includes English, and its formatting might not generalize to other EHRs. Third, as shown in Table LABEL:tab:prompt_eng, the specificity of criteria definitions has a large impact on the Assessment LLM’s ability to accurately assess patients. While this level of prompt engineering was still significantly easier than hand-engineering features with traditional NLP tools, and a real-world deployment would be able to leverage the much more detailed criteria definitions in a sponsor’s trial protocol, these results do still raise a generalization concern. Fourth, although our LLM-based system achieved state-of-the-art results on this benchmark, further evaluations are required to asses the safety and failure modes of applying these models in production settings. Fifth, the use of proprietary models such as GPT-4 may not be feasible for health systems that do not have HIPAA-compliant Azure instances. At a higher level, there are also open questions as to whether low enrollment in trials is a function of not being able to match patients (which this work addresses) or is instead primarily due to the stringency of the eligibility criteria themselves (Liu et al., 2021). While beyond the scope of this work, we believe that improving the design of eligibility criteria presents another interesting research direction which could benefit from the application of more sophisticated models such as LLMs (Liu et al., 2021).
@@ 7 Conclusion
Appendix A Compute Environment
Experiments are performed in a local on-prem university compute environment using 24 Intel Xeon 2.70GHz CPU cores, 8 Nvidia V100 GPUs, 4 Nvidia A100 GPUs, and 1.48 TB of RAM. All compute environments supported HIPAA-compliant data protocols.
Appendix B Dataset Selection
The n2c2 2018 cohort selection benchmark is simplified from real-world patient matching, and curating more realistic datasets is a key opportunity for the field (Stubbs et al., 2019). However, after considering multiple datasets for this paper, we selected the n2c2 2018 dataset for four primary reasons: (1) it is the largest publicly available trial matching dataset available in terms of both patients and depth of medical history available for each patient; (2) it is the only benchmark that contains full clinical notes rather than short ¡10 sentence case reports, and is thus the most ”realistic” for simulating the type of data present in a typical hospital’s EHR; (3) by virtue of being publicly available, it enables apples-to-apples comparisons with prior work and encourages reproducible science; (4) it provides criteria-level eligibility labels and not just trial-level labels, which allows for more granular evaluation. We only considered public benchmarks in order to enable fair comparisons with prior work (Stubbs et al., 2019). The other three main datasets we’ve seen evaluated in the literature (Jin et al., 2023; Nievas et al., 2023) are: TREC 2021 (Roberts et al., 2021), TREC 2022 (Roberts et al., 2021), and SIGIR 2016 (Koopman and Zuccon, 2016). Compared to the n2c2 2018 benchmark, however, these three datasets all have fewer patients (75, 50, and 60 respectively) and significantly less textual information associated with each patient (the patients are presented via ”case reports,” which are dense 5-10 sentence summaries of the patient’s entire relevant medical history written in plain English. This makes these datasets a poor fit for the task of clinical trial patient matching, as case reports are significantly different than the clinical text that would actually be encountered by a clinical trial matching system if it were deployed in a hospital setting. In contrast, the n2c2 dataset sources its text directly from deidentified clinical notes and thus offers a much more realistic distribution of text. Additionally, the brevity of case reports defeats the motivation for doing efficient retrieval and prompting (which was the focus of this work) – if you can easily fit everything in the context window of an LLM, then there is no need for the methods of this work. For example, the entire TREC 2021 dataset contains only 15,066 tokens of text, which means two entire copies of this dataset could fit into a single prompt of GPT-4, with room to spare (Roberts et al., 2021).
Appendix C Response Aggregation
When using the ACIN and ICIN prompting strategies, we get one eligibility prediction per criterion per note. This means a patient will have multiple eligibility decisions per criterion, and we therefore need a way to aggregate these decisions into one single decision per patient. To perform this aggregation, we use the following methodology on a criterion-by-criterion basis:
  1. 1.
    Criteria concerning the existence of a medical event across a patient’s entire medical history (e.g. having had an intra-abdominal surgery) are considered ”MET” if at least one note results in a ”MET” assessment by the LLM.
  2. 2.
    Criteria which are time-dependent (e.g. having taken a dietary supplement in the past 2 months) use the eligibility decision of the most recent note within the relevant time window.
The specific strategy used for each criterion is shown below.
Criteria Aggregation Method
ABDOMINAL max
ADVANCED-CAD max
ASP-FOR-MI max
CREATININE max
DRUG-ABUSE max
ALCOHOL-ABUSE max
HBA1C max
MAJOR-DIABETES max
MAKES-DECISIONS max
ENGLISH min
DIETSUPP-2MOS most recent (2 months)
KETO-1YR most recent (12 months)
MI-6MOS most recent (6 months)
Table S1: Criterion aggregation strategies
Figure S1: Criterion-level confusion matrices
Criteria True Value Prec. Rec. F1 Support
ABDOMINAL MET 0.7097 0.7333 0.7213 30
ABDOMINAL NOT MET 0.8545 0.8393 0.8468 56
ADVANCED-CAD MET 0.8542 0.9111 0.8817 45
ADVANCED-CAD NOT MET 0.8947 0.8293 0.8608 41
ALCOHOL-ABUSE MET 0.5000 0.6667 0.5714 3
ALCOHOL-ABUSE NOT MET 0.9878 0.9759 0.9818 83
ASP-FOR-MI MET 0.8590 0.9853 0.9178 68
ASP-FOR-MI NOT MET 0.8750 0.3889 0.5385 18
CREATININE MET 0.8333 0.8333 0.8333 24
CREATININE NOT MET 0.9355 0.9355 0.9355 62
DIETSUPP-2MOS MET 0.9286 0.8864 0.9070 44
DIETSUPP-2MOS NOT MET 0.8864 0.9286 0.9070 42
DRUG-ABUSE MET 0.7500 1.0000 0.8571 3
DRUG-ABUSE NOT MET 1.0000 0.9880 0.9939 83
ENGLISH MET 0.9863 0.9863 0.9863 73
ENGLISH NOT MET 0.9231 0.9231 0.9231 13
HBA1C MET 0.9394 0.8857 0.9118 35
HBA1C NOT MET 0.9245 0.9608 0.9423 51
KETO-1YR MET 0.0000 0.0000 0.0000 0
KETO-1YR NOT MET 1.0000 1.0000 1.0000 86
MAJOR-DIABETES MET 0.9730 0.8372 0.9000 43
MAJOR-DIABETES NOT MET 0.8571 0.9767 0.9130 43
MAKES-DECISIONS MET 0.9651 1.0000 0.9822 83
MAKES-DECISIONS NOT MET 0.0000 0.0000 0.0000 3
MI-6MOS MET 1.0000 0.7500 0.8571 8
MI-6MOS NOT MET 0.9750 1.0000 0.9873 78
Table S2: Criterion-level accuracies
Figure S2: Two-Stage Retrieval with $ACIN$ strategy
Model Prompt Strategy Change in Performance
Criteria Notes Prec. Rec. Overall Macro-F1 Overall Micro-F1
GPT-3.5 All All 0.04 0.03 0.05 0.03
GPT-3.5 All Individual 0.07 0.01 0.05 0.04
GPT-3.5 Individual All 0.05 0.03 0.05 0.03
GPT-3.5 Individual Individual 0.00 -0.01 0.01 0.01
GPT-4 All All -0.01 0.03 0.00 0.01
GPT-4 All Individual -0.04 0.01 -0.02 -0.02
GPT-4 Individual All -0.02 0.02 -0.01 -0.01
GPT-4 Individual Individual 0.00 0.02 0.01 -0.01
Table S3: Impact of Rationale field
Model Prompt Strategy Change in Performance
Criteria Notes Prec. Rec. Overall Macro-F1 Overall Micro-F1
GPT-4 All All -0.01 -0.02 -0.04 -0.02
GPT-4 All Individual -0.07 0.01 -0.05 -0.03
GPT-4 Individual All -0.01 -0.03 -0.06 -0.02
GPT-4 Individual Individual -0.01 0.02 -0.01 0.00
Table S4: Impact of 1-shot example
Criteria Definition Prevalence
ABDOMINAL
History of intra-abdominal surgery, small or large intestine resection, or small bowel obstruction
0.35
ADVANCED-CAD
Advanced cardiovascular disease (CAD). For the purposes of this annotation, we define “advanced” as having 2 or more of the following: • Taking 2 or more medications to treat CAD • History of myocardial infarction (MI) • Currently experiencing angina • Ischemia, past or present
0.52
ALCOHOL-ABUSE
Current alcohol use over weekly recommended limits
0.03
ASP-FOR-MI
Use of aspirin to prevent MI
0.79
CREATININE
Serum creatinine $>$ upper limit of normal
0.28
DIETSUPP-2MOS
Taken a dietary supplement (excluding vitamin D) in the past 2 months
0.51
DRUG-ABUSE
Drug abuse, current or past
0.03
ENGLISH
Patient must speak English
0.85
HBA1C
Any hemoglobin A1c (HbA1c) value between 6.5% and 9.5%
0.41
KETO-1YR
Diagnosis of ketoacidosis in the past year
0.00
MAJOR-DIABETES
Major diabetes-related complication. For the purposes of this annotation, we define “major complication” (as opposed to “minor complication”) as any of the following that are a result of (or strongly correlated with) uncontrolled diabetes: • Amputation • Kidney damage • Skin conditions • Retinopathy • nephropathy • neuropathy
0.50
MAKES-DECISIONS
Patient must make their own medical decisions
0.97
MI-6MOS
MI in the past 6 months
0.09
Table S5: 2018 n2c2 eligibility criteria description
Criteria Definition
ABDOMINAL
History of intra-abdominal surgery. This could include any form of intra-abdominal surgery, including but not limited to small/large intestine resection or small bowel obstruction
ADVANCED-CAD
Advanced cardiovascular disease (CAD). For the purposes of this annotation, we define “advanced” as having 2 or more of the following: (a) Taking 2 or more medications to treat CAD (b) History of myocardial infarction (MI) (c) Currently experiencing angina (d) Ischemia, past or present. The patient must have at least 2 of these categories (a,b,c,d) to meet this criterion, otherwise the patient does not meet this criterion. For ADVANCED-CAD, be strict in your evaluation of the patient – if they just have cardiovascular disease, then they do not meet this criterion.
ALCOHOL-ABUSE
Current alcohol use over weekly recommended limits
ASP-FOR-MI
Use of aspirin for preventing myocardial infarction (MI).
CREATININE
Serum creatinine level above the upper normal limit
DIETSUPP-2MOS
Consumption of a dietary supplement (excluding vitamin D) in the past 2 months. To assess this criterion, go through the list of medications_and_supplements taken from the note. If a substance could potentially be used as a dietary supplement (i.e. it is commonly used as a dietary supplement, even if it is not explicitly stated as being used as a dietary supplement), then the patient meets this criterion. Be lenient and broad in what is considered a dietary supplement. For example, a ’multivitamin’ and ’calcium carbonate’ should always be considered a dietary supplement if they are included in this list.
DRUG-ABUSE
Current or past history of drug abuse
ENGLISH
Patient speaks English. Assume that the patient speaks English, unless otherwise explicitly noted. If the patientś language is not mentioned in the note, then assume they speak English and thus meet this criteria.
HBA1C
Any hemoglobin A1c (HbA1c) value between 6.5% and 9.5%
KETO-1YR
Diagnosis of ketoacidosis within the past year
MAJOR-DIABETES
Major diabetes-related complication. Examples of “major complication” (as opposed to “minor complication”) include, but are not limited to, any of the following that are a result of (or strongly correlated with) uncontrolled diabetes: • Amputation • Kidney damage • Skin conditions • Retinopathy • nephropathy • neuropathy. Additionally, if multiple conditions together imply a severe case of diabetes, then count that as a major complication.
MAKES-DECISIONS
Patient must make their own medical decisions. Assume that the patient makes their own medical decisions, unless otherwise explicitly noted. There is no information provided about the patientś ability to make their own medical decisions, then assume they do make their own decisions and therefore meet this criteria
MI-6MOS
Myocardial infarction (MI) within the past 6 months
Table S6: Redefined criteria definitions
Figure S3: Prompt used for All Criteria setting
Figure S4: Prompt used for Individual Criteria setting